# UV-Attack: Physical-World Adversarial Attacks for Person Detection via Dynamic-NeRF-based UV Mapping

### v1

@@ Abstract
In recent research, adversarial attacks on person detectors using patches or static 3D model-based texture modifications have struggled with low success rates due to the flexible nature of human movement. Modeling the 3D deformations caused by various actions has been a major challenge. Fortunately, advancements in Neural Radiance Fields (NeRF) for dynamic human modeling offer new possibilities. In this paper, we introduce UV-Attack, a groundbreaking approach that achieves high success rates even with extensive and unseen human actions. We address the challenge above by leveraging dynamic-NeRF-based UV mapping. UV-Attack can generate human images across diverse actions and viewpoints, and even create novel actions by sampling from the SMPL parameter space. While dynamic NeRF models are capable of modeling human bodies, modifying clothing textures is challenging because they are embedded in neural network parameters. To tackle this, UV-Attack generates UV maps instead of RGB images and modifies the texture stacks. This approach enables real-time texture edits and makes the attack more practical. We also propose a novel Expectation over Pose Transformation loss (EoPT) to improve the evasion success rate on unseen poses and views. Our experiments show that UV-Attack achieves a 92.75% attack success rate against the FastRCNN model across varied poses in dynamic video settings, significantly outperforming the state-of-the-art AdvCamou attack, which only had a 28.50% ASR. Moreover, we achieve 49.5% ASR on the latest YOLOv8 detector in black-box settings. This work highlights the potential of dynamic NeRF-based UV mapping for creating more effective adversarial attacks on person detectors, addressing key challenges in modeling human movement and texture modification.
@@ 1 Introduction
We compare UV-Attack with the latest person detection attacks (Hu et al., 2022; 2023) and diffusion-based attacks (Xue et al., 2023; Lin et al., 2023). Our experiments demonstrate the superiority of our approach over in terms of evasion success rates in free-pose scenarios. As shown in Figure 1, we physically printed the Bohemian-style clothing and successfully fooled person detectors in dynamic video settings, even when the subjects were in continuous and significant movement. Our key contributions are:
@@
    We propose a novel person detection attack algorithm that effectively handles pose variation through UV mapping. Our method can generate unseen actions and can edit human clothes in real-time. Moreover, we propose a novel Expectation over Pose Transformation loss (EoPT) to improve the evasion success rate on unseen poses and views.
@@
    Figure 1: Visualization of different person detector attacks. We compare our attack with different person detection attacks, including Adv-Yolo (Thys et al., 2019), Adv-Tshirt (Xu et al., 2020),NatualPatch (Hu et al., 2021), Adv-Texture (Hu et al., 2022) and Adv-Camou (Hu et al., 2023). Our attacks achieve higher ASR under diverse poses and viewpoints than previous attacks.
@@ 3.1 Preliminary: Dynamic Neural Radiance Fields
NeRF enables novel-view photorealistic synthesis by modeling 3D objects as a continuous volumetric radiance field allowing control over illuminations and materials (Mildenhall et al., 2021). It can be formulated as $F:(\mathbf{x},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$, where $F$ is a MLP and takes the 3D coordinate $\mathbf{x}\in\mathbb{R}^{3}$ and the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as input and outputs the corresponding color c and volume density $\sigma$. Then NeRF computes the pixel color $C(\textbf{r})$ by accumulating all colors along a ray $\textbf{r}=\text{o}+t\textbf{d}$ that emitting from the camera’s original point $o$ and passing the pixel with near and far bounds $t_{n}$ and $t_{f}$,
[…]
@@
The core idea of UV-Volume is to render a NeRF model into UV maps rather than RGB images. UV maps contain UV coordinates (or texture coordinates), which are assigned pixel coordinates of a predefined texture image $\mathcal{E}$. The UV mapping process assigns pixels in the predefined texture image $\mathcal{E}$ to a rendered RGB image according to the UV coordinates. The UV-Volume input is the pose parameter $\mathbf{\theta}$, the camera position $\eta$ and the light parameter $\tau$. It firstly computes a pixel feature $\mathcal{F}(\mathbf{r})$ by a rendering equation similar to Eq. 1, where r is a camera ray. Then $\mathcal{F}(\bm{r})$ is fed into a UV decoder and is decoded as part assignment and UV coordinates through $[\hat{\mathcal{P}}(\mathbf{r}),\hat{\mathcal{U}}(\mathbf{r}),\hat{\mathcal{V}}(\mathbf{r})]=M_{uv}(\mathcal{F}(\mathbf{r}))$. $\hat{\mathcal{P}}$ is a part assignment corresponding to different human body parts.
UV-volume also incorporates a texture generator, denoted by $G$, to generate texture stack images $\mathcal{E}_{k}=G(\theta,\bm{k})$, where $\bm{k}$ is a one-hot vector that specifies the body part (e.g., arm, leg). A texture feature, $\mathbf{e}_{k}$, is then sampled from $\mathcal{E}_{k}$ based on the UV coordinate using the differentiable grid sampling: $\mathbf{e}_{k}(\mathbf{r})=\mathcal{E}_{k}[\mathcal{\hat{U}}_{k}(\mathbf{r}),\mathcal{\hat{V}}_{k}(\mathbf{r})]$. Finally, the RGB color for the camera ray $\bm{r}$ is calculated as a weighted sum of individual body part colors $\hat{C}_{k}(\mathbf{r})$: $\hat{C}(\mathbf{r})=\Sigma_{k=1}^{K}\hat{\mathcal{P}}_{k}(\mathbf{r})\hat{C}_{k}(\mathbf{r}).$ Each $\hat{C}_{k}(\mathbf{r})$ is determined by a MLP, $M_{c}$, which takes as input the UV coordinates (with positional encoding), the sampled texture feature, and the encoded view direction:
@@
where $K$ is the total number of human body parts, $\gamma(\cdot)$ is a positional encoding function and $M_{c}$ is an MLP to map UV coordinates, sampled texture pixels and view directions into RGB images.
Figure 2: The pipeline of UV-Attack. UV-Attack first samples random pose parameters from a Gaussian Mixture Model. We then use the sampled pose, camera and light parameters to generate IUV maps and texture stacks. We employ a pre-trained stable diffusion model and modify the initial latent to generate adversarial patches. Finally, we modify the texture stacks using the adversarial patch and get the human image in RGB space. The solid lines below indicate the direction of gradient propagation, while the dashed lines above do not require gradients.
@@ 3.2 The pipeline of UV-Attack
To generate unseen poses for the target person, we use a Gaussian Mixture Model (GMM) to fit a human pose distribution on target person images and an annotated pose dataset MPII (Andriluka et al., 2014). Then we sample pose parameters from this Gaussian Mixture distribution as the UV-Volume model’s input. Specifically, We first use the SPIN model (Kolotouros et al., 2019) to estimate the poses $\hat{\bm{\theta}}$ and shapes parameters $\hat{\bm{\beta}}$ from the target person’s video and add auxiliary poses $\tilde{\bm{\theta}}$ from public annotated poses datasets like MPII (Andriluka et al., 2014). Then we parameterize the distribution $p([\hat{\bm{\theta}},\tilde{\bm{\theta}}])$ by a mixture of $K$ Gaussian components and take the transformation of random variable approach to ensure the support of $p([\hat{\bm{\theta}},\tilde{\bm{\theta}}])$ is in scope of $[\bm{\theta}_{min},\bm{\theta}_{max}]$ by
@@
To improve the speed of adversarial attacks, it’s crucial to efficiently modify and render clothing on 3D human models. We slightly adjust the UV-Volume architecture to further improve the texture modification speed. Specifically. we firstly define linearly spaced vectors $\mathbf{u},\mathbf{v}\in\text{linespace}(0,1,N_{t})$, where $N_{t}$ is the texture’s size. Then we create 2D grids using a mesh grid defined as $[\tilde{\mathcal{U}},\tilde{\mathcal{V}}]=[\mathbf{1}_{N_{t}}\otimes\mathbf{u},(\mathbf{1}_{N_{t}}\otimes\mathbf{v})^{T}]$, where is $\mathbf{1}_{N_{t}}$ is a column vector of ones with dimension $N_{t}$ and $\otimes$ denotes the Kronecker product. Then we remove the viewing direction $\bm{d}$ from Eq.2 to generate the view-independent texture stacks Besides the texture stack $\tilde{C}$, we also generate IUV maps according to the pose, camera, and light parameters, which can be expressed $x_{\text{IUV}}=[\hat{\mathcal{P}},\hat{\mathcal{U}},\hat{\mathcal{V}}]=M_{uv}(\mathcal{F}_{\theta,\eta,\tau})$, where $\theta$ is the pose parameter and $\eta,\tau$ are the camera and light parameters.
During the process of image rendering for new actions, we randomly sample the camera views from different elevation and azimuth angles and different light positions and backgrounds. The rendered images can be expressed as
@@
where $x_{adv}$ is an adversarial patch generated by the latent diffusion model, $\tau$ is the camera and light parameters and $M$ a human mask that can be estimated by the accumulated transmittance $M=\exp\left(-\sum_{j=1}^{N_{i}-1}\sigma(\mathbf{x}_{j})\delta_{j}\right)$. We simply repeat the adversarial patch on the texture stacks, which have shown to be able to improve the attack robustness (Hu et al., 2022). Because of the texture stacks are disentangled, we can easily modify any part of the human body. Then we distort the texture stacks through Thin Plate Spline (TPS) (Hu et al., 2023) to simulate cloth distortions. Finally, we use the grid sampling function $g(\cdot,\cdot)$ to sample pixel colors from the modified texture stack according to the IUV map $x_{\text{IUV}}$ and combine it with background image $x_{b}$.
@@
Diffusion models have been found to be able to enhance the transferability of adversarial examples (Xue et al., 2023; Lin et al., 2023; Chen et al., 2023b). However, previous work either modifying the intermediate latent (Chen et al., 2023a; b) or diffusing and denoising an initial image (Lin et al., 2023; Xue et al., 2023), which has limited noise levels, resulting in impractical for physical adversarial attacks. UV-Attack contaminates the initial, highly noisy latent variable $z_{T}$ of a pre-trained stable diffusion model (Rombach et al., 2022) and remove classifier-free guidance. These approaches facilitate more substantial modifications and explore broader latent spaces and less familiar areas, resulting in significantly higher physical attack success rates while maintaining the transferability. However, due to the larger optimization space, relying solely on gradient-based algorithms may lead to local minima. To enhance optimization efficiency, we combined black-box and white-box algorithms. Initially, we employed the Particle Swarm Optimization (PSO) algorithm to find a suitable starting point. Subsequently, we utilized the Adam optimizer to fine-tune this point to further decrease the loss. To exclude the influence of stable diffusion prompts, we evaluated the average ASR by using different class labels except "person" in the COCO dataset as prompts.
@@
6:     Evaluate fitness ($-\mathcal{L}_{adv}$) of particles and update $p_{i}^{(t)}$ and $g^{(t)}$
@@ 4.1 Experimental Settings
We evaluate the success rate of digital attacks on the ZJU-Mocap dataset (Peng et al., 2021). For the physical attack, we collected videos of five individuals and used SPIN and DensePose to extract SMPL parameters and pseudo-supervised IUV maps for training the UV-Volumes. The model training and attack implementation are conducted on a single Nvidia 3090 GPU. We utilize a pretrained stable-diffusion model finetuned on 256$\times$256 images. The diffusion process was set to 10 steps. The Particle Swarm Optimization (PSO) ran for 30 epochs with 50 swarms, and the Adam optimizer ran for 300 epochs. We collected 100 different backgrounds from both indoor and outdoor scenarios. For each epoch, we randomly sampled 100 poses, camera and light positions, and backgrounds for training. The camera is sampled from $\text{azim}\in[-180^{\circ},180^{\circ}]$ and $\text{elev}\in[0^{\circ},30^{\circ}]$.
@@ 4.2 Digital Attack Results
(A) Dataset and baseline settings. To evaluate the ASR of different attacks under unseen pose scenarios, we randomly sample 1000 poses from the GMM and combine them with different backgrounds and camera and light conditions to construct an unseen pose dataset. Then we evaluate the average ASR on five different individuals from the ZJU-Mocap dataset, as shown in Figure 6. Because previous attacks did not consider the multi-pose scenarios, they cannot be directly evaluated. Therefore, we combine their methodologies with the UV-Volume model to generate images under different poses for test. For AdvCamou (Hu et al., 2023), we first train a Voronoi diagram using their methods and then apply it to the texture stacks under a fixed pose. Then we test the generated Voronoi diagram on the multi-pose dataset. For other patch-based attacks, we use a fixed pose and put the generated patch on the texture stacks for training, followed by testing on the multi-pose dataset. Additionally, we added two additional baselines. One is to use the latent diffusion model with fixed poses to train the adversarial textures, referred to as the LDM attack. The other is to sample frames from a multi-pose video and then extract the pose parameters for the UV-Volume, referred to as $\text{Ours}_{\text{video}}$. (B) Results analysis. We present the test ASRs under different IoU thresholds in Table 1. All attacks were trained and tested on the same detectors . It indicates that our attack significantly outperforms previous attacks under unseen pose scenarios. Additionally, our attack maintains consistently high ASRs as the IoU thresholds decrease, in contrast to some previous attacks like AdvTexture which significantly drops in performance. Experimental results demonstrate that our method can significantly improve attack resilience against pose variations. In addition, it is shown that sampling pose parameters from a GMM distribution rather than from video frames can improve the ASR. This is because we can synthesize unseen poses for training.
@@
(A) We trained our attacks on FastRCNN and YOLOv3, respectively, and tested the transferability on various person detectors. Because AdvCamou is based on a static 3D model, we disable pose sampling in the training and testing in this table for fair comparisons. As shown in Table 2, our attack trained on FastRCNN has high transferability on Deformable-DETR, MaskRCNN, and RetinaNet. It is also shown that previous diffusion-based attacks, Diff2Conf and DiffPGD, have very low ASRs in our attack settings, even on the white-box models. We think this is because only adding noise to an initial image and then denoising it can not fully explore the LDM’s latent space and, therefore cannot find an optimum latent. In addition, our attack achieves high ASRs on the latest YOLOv8 detector (Jocher et al., 2023), while previous attacks did not perform very well on it.

### v2

# UV-Attack: Physical-World Adversarial Attacks on Person Detection via Dynamic-NeRF-based UV Mapping

@@ Abstract
In recent research, adversarial attacks on person detectors using patches or static 3D model-based texture modifications have struggled with low success rates due to the flexible nature of human movement. Modeling the 3D deformations caused by various actions has been a major challenge. Fortunately, advancements in Neural Radiance Fields (NeRF) for dynamic human modeling offer new possibilities. In this paper, we introduce UV-Attack, a groundbreaking approach that achieves high success rates even with extensive and unseen human actions. We address the challenge above by leveraging dynamic-NeRF-based UV mapping. UV-Attack can generate human images across diverse actions and viewpoints, and even create novel actions by sampling from the SMPL parameter space. While dynamic NeRF models are capable of modeling human bodies, modifying clothing textures is challenging because they are embedded in neural network parameters. To tackle this, UV-Attack generates UV maps instead of RGB images and modifies the texture stacks. This approach enables real-time texture edits and makes the attack more practical. We also propose a novel Expectation over Pose Transformation loss (EoPT) to improve the evasion success rate on unseen poses and views. Our experiments show that UV-Attack achieves a 92.7% attack success rate against the FastRCNN model across varied poses in dynamic video settings, significantly outperforming the state-of-the-art AdvCamou attack, which only had a 28.5% ASR. Moreover, we achieve 49.5% ASR on the latest YOLOv8 detector in black-box settings. This work highlights the potential of dynamic NeRF-based UV mapping for creating more effective adversarial attacks on person detectors, addressing key challenges in modeling human movement and texture modification. The code is available at https://github.com/PolyLiYJ/UV-Attack.
@@ 1 Introduction
We compare UV-Attack with the latest person detection attacks (Hu et al., 2022; 2023) and diffusion-based attacks (Xue et al., 2023; Lin et al., 2023). Our experiments demonstrate the superiority of our approach in terms of evasion success rates in free-pose scenarios. As shown in Figure 1, we physically printed the Bohemian-style clothing and successfully fooled person detectors in dynamic video settings, even when the subjects were in continuous and significant movement. Our key contributions are:
@@
    We propose a novel person detection attack algorithm that effectively handles pose variation through UV mapping. Our method can generate unseen actions and can edit human clothes in real time. Moreover, we propose a novel Expectation over Pose Transformation loss (EoPT) to improve the evasion success rate on unseen poses and views.
@@
    Figure 1: Visualization of different person detector attacks. We compare our attack with different person detection attacks, including Adv-Yolo (Thys et al., 2019), Adv-Tshirt (Xu et al., 2020), NaturalPatch (Hu et al., 2021), Adv-Texture (Hu et al., 2022) and Adv-Camou (Hu et al., 2023). Our attacks achieve higher ASR under diverse poses and viewpoints than previous attacks.
@@ 3.1 Preliminary: Dynamic Neural Radiance Fields
NeRF enables novel-view photorealistic synthesis by modeling 3D objects as a continuous volumetric radiance field, allowing control over illuminations and materials (Mildenhall et al., 2021). It can be formulated as $F:(\mathbf{x},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$, where $F$ is a MLP and takes the 3D coordinate $\mathbf{x}\in\mathbb{R}^{3}$ and the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as input and outputs the corresponding color c and volume density $\sigma$. Then NeRF computes the pixel color $C(\textbf{r})$ by accumulating all colors along a ray $\textbf{r}=\text{o}+t\textbf{d}$ that emitting from the camera’s original point $o$ and passing the pixel with near and far bounds $t_{n}$ and $t_{f}$,
$C(\textbf{r})=\int_{t_{n}}^{t_{f}}T(t)\cdot\sigma(\mathbf{r}(t))\cdot c(\mathbf{r}(t),\mathbf{d})\,dt,\;\;\text{where}\;T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\,ds\right),$ (1)
@@
The core idea of UV-Volume is to render a NeRF model into UV maps rather than RGB images. UV maps contain UV coordinates (or texture coordinates), which are assigned pixel coordinates of a predefined texture image $\mathcal{E}$. The UV mapping process assigns pixels in the predefined texture image $\mathcal{E}$ to a rendered RGB image according to the UV coordinates. The UV-Volume input is the pose parameter $\mathbf{\theta}$, the camera position $\eta$ and the light parameter $\tau$. It firstly computes a pixel feature $\mathcal{F}(\mathbf{r})$ by a rendering equation similar to Eq. 1, where r is a camera ray. Then $\mathcal{F}(\bm{r})$ is fed into a UV decoder and is decoded as part assignment and UV coordinates through $\hat{\mathcal{P}}(\mathbf{r}),\hat{\mathcal{U}}(\mathbf{r}),\hat{\mathcal{V}}(\mathbf{r})]=M_{uv}(\mathcal{F}(\mathbf{r}))$. $\hat{\mathcal{P}}$ is a part assignment corresponding to different human body parts.
UV-volume also incorporates a texture generator, denoted by $G$, to generate texture stack images $\mathcal{E}_{k}=G(\theta,\bm{k})$, where $\bm{k}$ is a one-hot vector that specifies the body part (e.g., arm, leg). A texture feature, $\mathbf{e}_{k}$, is then sampled from $\mathcal{E}_{k}$ based on the UV coordinate using the differentiable grid sampling: $\mathbf{e}_{k}(\mathbf{r})=\mathcal{E}_{k}[\mathcal{\hat{U}}_{k}(\mathbf{r}),\mathcal{\hat{V}}_{k}(\mathbf{r})]$. Finally, the RGB color for the camera ray $\bm{r}$ is calculated as a weighted sum of individual body part colors $\hat{C}_{k}(\mathbf{r})$: $\hat{C}(\mathbf{r})=\Sigma_{k=1}^{K}\hat{\mathcal{P}}_{k}(\mathbf{r})\hat{C}_{k}(\mathbf{r}).$ Each $\hat{C}_{k}(\mathbf{r})$ is determined by an MLP, $M_{c}$, which takes as input the UV coordinates (with positional encoding), the sampled texture feature, and the encoded view direction:
@@
where $K$ is the total number of human body parts and $\gamma(\cdot)$ a positional encoding function. The $M_{c}$ is an MLP that maps UV coordinates, sampled texture pixels and view directions into RGB images.
Figure 2: The pipeline of UV-Attack. UV-Attack first samples random pose parameters from a Gaussian Mixture Model. We then use the sampled pose, camera, and light parameters to generate IUV maps and texture stacks. We employ a pre-trained stable diffusion model and modify the initial latent to generate adversarial patches. Finally, we modify the texture stacks using the adversarial patch and get the human image in RGB space. The solid lines below indicate the direction of gradient propagation, while the dashed lines above do not require gradients.
@@ 3.2 The pipeline of UV-Attack
[…]
@@
To improve the speed of adversarial attacks, it’s crucial to efficiently modify and render clothing on 3D human models. We slightly adjust the UV-Volume architecture to further improve the texture modification speed. Specifically. we firstly define linearly spaced vectors $\mathbf{u},\mathbf{v}\in\text{linespace}(0,1,N_{t})$, where $N_{t}$ is the texture’s size. Then we create 2D grids using a mesh grid defined as $\tilde{\mathcal{U}},\tilde{\mathcal{V}}]=[\mathbf{1}_{N_{t}}\otimes\mathbf{u},(\mathbf{1}_{N_{t}}\otimes\mathbf{v})^{T}]$, where is $\mathbf{1}_{N_{t}}$ is a column vector of ones with dimension $N_{t}$ and $\otimes$ denotes the Kronecker product. Then we remove the viewing direction $\bm{d}$ from Eq.2 to generate the view-independent texture stacks Besides the texture stack $\tilde{C}$, we also generate IUV maps according to the pose, camera, and light parameters, which can be expressed $x_{\text{IUV}}=[\hat{\mathcal{P}},\hat{\mathcal{U}},\hat{\mathcal{V}}]=M_{uv}(\mathcal{F}_{\theta,\eta,\tau})$, where $\theta$ is the pose parameter and $\eta,\tau$ are the camera and light parameters.
During the image rendering process for new actions, we randomly sample the camera views from different elevation and azimuth angles and different light positions and backgrounds. The rendered images can be expressed as
@@
where $x_{adv}$ is an adversarial patch generated by the latent diffusion model, $\tau$ is the camera and light parameters and $M$ a human mask that can be estimated by the accumulated transmittance $M=\exp\left(-\sum_{j=1}^{N_{i}-1}\sigma(\mathbf{x}_{j})\delta_{j}\right)$. We simply repeat the adversarial patch on the texture stacks, which have shown to be able to improve the attack robustness (Hu et al., 2022). Because the texture stacks are disentangled, we can easily modify any part of the human body. Then we distort the texture stacks through Thin Plate Spline (TPS) (Hu et al., 2023) to simulate cloth distortions. Finally, we use the grid sampling function $g(\cdot,\cdot)$ to sample pixel colors from the modified texture stack according to the IUV map $x_{\text{IUV}}$ and combine it with background image $x_{b}$.
@@
Diffusion models have been found to be able to enhance the transferability of adversarial examples (Xue et al., 2023; Lin et al., 2023; Chen et al., 2023b). However, previous work either modifying the intermediate latent (Chen et al., 2023a; b) or diffusing and denoising an initial image (Lin et al., 2023; Xue et al., 2023), which has limited noise levels, resulting in impractical for physical adversarial attacks. UV-Attack contaminates the initial, highly noisy latent variable $z_{T}$ of a pre-trained stable diffusion model (Rombach et al., 2022) and remove classifier-free guidance. These approaches facilitate more substantial modifications and explore broader latent spaces and less familiar areas, resulting in significantly higher physical attack success rates while maintaining transferability. However, due to the larger optimization space, relying solely on gradient-based algorithms may lead to local minima. To enhance optimization efficiency, we combined black-box and white-box algorithms. Initially, we employed the Particle Swarm Optimization (PSO) algorithm to find a suitable starting point. Subsequently, we utilized the Adam optimizer to fine-tune this point to further decrease the loss. To exclude the influence of stable diffusion prompts, we evaluated the average ASR by using different class labels except "person" in the COCO dataset as prompts.
@@
6:  Evaluate fitness ($\mathcal{L}_{adv}$) of particles and update $p_{i}^{(t)}$ and $g^{(t)}$
@@ 4.1 Experimental Settings
We evaluate the success rate of digital attacks on the ZJU-Mocap dataset (Peng et al., 2021). For the physical attack, we collected videos of five individuals and used SPIN and DensePose to extract SMPL parameters and pseudo-supervised IUV maps for training the UV-Volumes. The model training and attack implementation are conducted on a single Nvidia 3090 GPU. We utilize a pretrained stable-diffusion model finetuned on $\times256$ images. The diffusion process was set to 10 steps. The Particle Swarm Optimization (PSO) ran for 30 epochs with 50 swarms, and the Adam optimizer ran for 300 epochs. We collected 100 different backgrounds from both indoor and outdoor scenarios. For each epoch, we randomly sampled 100 poses, camera and light positions, and backgrounds for training. The camera is sampled from $\text{azim}\in[-180^{\circ},180^{\circ}]$ and $\text{elev}\in[0^{\circ},30^{\circ}]$.
@@ 4.2 Digital Attack Results
(A) Dataset and baseline settings. To evaluate the ASR of different attacks under unseen pose scenarios, we randomly sample 1000 poses from the GMM and combine them with different backgrounds and camera and light conditions to construct an unseen pose dataset. Then we evaluate the average ASR on five different individuals from the ZJU-Mocap dataset, as shown in Figure 6. Because previous attacks did not consider the multi-pose scenarios, they cannot be directly evaluated. Therefore, we combine their methodologies with the UV-Volume model to generate images under different poses for testing. For AdvCamou (Hu et al., 2023), we first train a Voronoi diagram using their methods and then apply it to the texture stacks under a fixed pose. Then we test the generated Voronoi diagram on the multi-pose dataset. For other patch-based attacks, we use a fixed pose and put the generated patch on the texture stacks for training, followed by testing on the multi-pose dataset. Additionally, we added two additional baselines. One is to use the latent diffusion model with fixed poses to train the adversarial textures, referred to as the LDM attack. The other is to sample frames from a multi-pose video and then extract the pose parameters for the UV-Volume, referred to as $\text{Ours}_{\text{video}}$. (B) Results analysis. We present the test ASRs under different IoU thresholds in Table 1. All attacks were trained and tested on the same detectors . It indicates that our attack significantly outperforms previous attacks under unseen pose scenarios. Additionally, our attack maintains consistently high ASRs as the IoU thresholds decrease, in contrast to some previous attacks like AdvTexture, which significantly drops in performance. Experimental results demonstrate that our method can significantly improve attack resilience against pose variations. In addition, it is shown that sampling pose parameters from a GMM distribution rather than from video frames can improve the ASR. This is because we can synthesize unseen poses for training.
@@
(A) We trained our attacks on FastRCNN and YOLOv3, respectively, and tested the transferability on various person detectors. Because AdvCamou is based on a static 3D model, we disable pose sampling in the training and testing in this table for fair comparisons. As shown in Table 2, our attack trained on FastRCNN has high transferability on Deformable-DETR, MaskRCNN, and RetinaNet. It is also shown that previous diffusion-based attacks, Diff2Conf and DiffPGD, have very low ASRs in our attack settings, even on the white-box models. We think this is because only adding noise to an initial image and then denoising it can not fully explore the LDM’s latent space and, therefore, cannot find an optimum latent. In addition, our attack achieves high ASRs on the latest YOLOv8 detector (Jocher et al., 2023), while previous attacks did not perform very well on it.